# Who is Responsible? Data, Models, or Regulations, A Comprehensive Survey on Responsible Generative AI for a Sustainable Future

### v6

@@ 6.1. Synthesis and Interpretation
Our review shows that post-GPT scholarship concentrates on five areas: safety and benchmarking, fairness and governance, explainability and transparency, privacy and security, and sectoral applications. Bias and toxicity receive dense coverage, while deepfakes, sustainability, and system-level failure modes receive relatively less attention. This pattern aligns with the thematic trends we documented and with the growing prominence of multimodal and agentic systems in 2022–2025, which increases the need for evaluation that captures tool use and context.
Closing the accountability chain
@@
To fully answer our primary research question, Who is Responsible?, our synthesis posits that the user must be transitioned from a passive consumer to a systemic guardrail. The current imbalance in GenAI governance, where 74% of organizations use GenAI but only 26% have a formal strategy, often results in the end-user inheriting unmitigated technical risks like PII leakage and hallucination. By triangulating the sectoral failures in Section 5 with the digital literacy gaps identified in Section 3.2.2, we propose a Symmetric Responsibility Model. In this model, developers are accountable for Transparency and Alignment, while users are accountable for Operational Integrity and Dissemination. Digital literacy, therefore, is not merely a skill but a core technical control that must be assessed alongside model-side benchmarks in Section 4.2. Without this user centric governance, even perfectly aligned models remain vulnerable to the last mile failure of human misinterpretation or malicious re contextualization. Table 10 operationalizes this model by mapping specific failure modes to the reciprocal duties of each stakeholder.
Table 10. The Symmetric Responsibility Matrix: Mapping Failure Modes to Stakeholder Duties
Failure Mode Developer Duty (Engineering Controls) User Duty (Operational Integrity)
Hallucination & Misinfo Implement RAG frameworks and calibrated uncertainty reporting (KPI 9) (Wu et al., 2024). Verify factual claims via primary sources and cross modal grounding checks.
PII & Data Leakage Deploy Differential Privacy and automated scrubbing during inference (KPI 2) (Nasr et al., 2023). Practice prompt hygiene by omitting sensitive or proprietary data from inputs.
Adversarial Jailbreaking Conduct continuous red teaming and implement robust system guardrails (C8) (Samvelyan et al., 2024). Adhere to acceptable use policies and report emergent safety bypasses to developers.
Representational Bias Perform fairness audits using intersectional metrics like SPD and DI (KPIs 3 to 4) (Kun, 2015). Critically appraise outputs for stereotyping and contextualize results within local norms.
Over reliance (Automation Bias) Provide model confidence scores and explicit explanation gap warnings (KPI 8) (Strathern, 1997a). Maintain human in the loop (HITL) oversight and preserve independent domain expertise.
Industry Safety Artifacts and Advanced Interpretability.
@@
While academic benchmarks offer standardized metrics, the operational reality of ResGenAI is increasingly defined by safety documentation from frontier labs and emerging mechanistic interpretability techniques, as synthesized in Figure 11. Frontier labs have established documentation protocols as de facto benchmarks for governance facing evidence, directly mapping to our C1 to C10 rubric. OpenAI system cards for GPT 4 and GPT 4o provide granular audits for sycophancy and PII leakage, demonstrating high maturity applications of metric validity through traceable safety refusal evidence. Anthropic utilizes Constitutional AI to automate alignment via RLAIF, outlining tiered safeguards that address malicious actor realism. Meta Llama 3 documentation introduces standardized protocols for multilingual red teaming, vital for addressing cross cultural ethical gaps. Google DeepMind technical reports provide empirical evidence for bias and accuracy trade offs, detailing the overcorrection mechanisms that led to observed representational failures. This industry evidence is complemented by the shift toward mechanistic interpretability, which maps internal neural activations to semantic concepts to offer a more rigorous safety check than behavior level prompts alone. Prominent techniques like Sparse Autoencoders (SAEs) isolate circuits responsible for deceptive or biased reasoning, enabling model steering to suppress harmful features during inference. Integrating these mechanistic audits into the RAI lifecycle addresses the black box challenge by enabling empirical verification of internal safety alignment before deployment.
@@
Figure 11. Integration of industry safety documentation and advanced mechanistic interpretability for governance facing evidence.
ResGenAI sustainability via lifecycle assessment
@@
Sustainability in ResGenAI must transition from qualitative discourse to rigorous quantification grounded in Lifecycle Assessment (LCA) frameworks, as integrated into our governance landscape (Figure 8) and synthesized in Figure 12. This necessitates formalizing carbon reporting and energy benchmarks across the development cycle, including pretraining, fine tuning, and inference. Following our proposed lifecycle KPIs (Section 4.2.4), we advocate for reporting energy consumption and carbon equivalent emissions normalized by throughput, such as per 1K tokens generated. Seminal research indicates that training large scale generative models produces environmental footprints comparable to the lifetime emissions of multiple automobiles, highlighting the need for transparent tracking tools and hardware aware reporting. A sustainable AI ecosystem requires a dual focus on optimizing the computational efficiency of models and leveraging generative insights to advance ecological goals, ensuring that technical innovation supports long term environmental stability.
Figure 12. Quantifying Responsible GenAI Sustainability via Lifecycle Assessment (LCA) in ResGenAI Governance. This framework formalizes carbon reporting and energy benchmarks across the pretraining, fine tuning, and inference stages, enabling a transition to rigorous quantification and supporting a sustainable AI ecosystem.
6.2. Lessons from Industry Deployment Incidents
While our benchmark analysis identifies risk-coverage gaps analytically, real-world deployment incidents from major AI developers offer empirical grounding. We synthesize prominent cases below, mapped to our rubric criteria, to demonstrate how the gaps in Tables 36 manifest in production systems.
@@ 6.2. Lessons from Industry Deployment Incidents
Alignment and safety failures (C2, C3, C8) Microsoft Bing Chat integration of GPT-4 in early 2023 produced adversarial persona adoption and emotional manipulation during extended sessions (Roose, 2023), behaviors that emerged through multi-turn prompting, precisely the adaptive adversary pattern captured by C8, for which only 2 of 21 benchmarks scored above zero. OpenAI GPT-4o system card documented sycophantic and manipulative voice-mode outputs 444developers.openai.com/api/docs/models/gpt-4o, illustrating subtler alignment failures beyond toxicity classifiers. Google’s Gemini image generator was suspended in February 2024 after producing historically inaccurate racially diverse depictions of historical figures555blog.google/products/gemini/gemini-image-generation-issue/, demonstrating how overcorrection in bias mitigation (C1) can itself produce fairness failures, an interaction that benchmarks evaluate in isolation.
@@
Hallucination in deployed systems (C4, C9) In Mata v. Avianca (2023), attorneys submitted a brief with six fabricated ChatGPT-generated case citations, resulting in judicial sanctions666caselaw.findlaw.com/court/us-dis-crt-sd-new-yor/2335142.html. These cases underscore that only 7 of 21 benchmarks address misinformation with documented constructs, and that the high-stakes error rate KPI ($E_{h}$, Section 4.2.4) must be operationalized in deployment settings.
@@
Privacy and data leakage (C6) Samsung banned internal ChatGPT use in 2023 after employees inadvertently leaked proprietary source code through conversations (Samsung Electronics, 2023), a usage-path leakage vector that no current privacy benchmark assesses. Nasr et al. (Nasr et al., 2023) demonstrated scalable extraction of memorized training data, including PII, from production ChatGPT via divergence attacks, exceeding the scope of PrivLM-Bench and PrivacyLens.
@@
Deepfake and media integrity (C5) A Hong Kong finance worker transferred approximately \$25 million after a video call with deepfake recreations of company executives (Chen, 2024), and AI-generated robocalls impersonating President Biden targeted the 2024 New Hampshire primary (New Hampshire Department of Justice, 2024). Only SHIELD provides explicit C5 evaluation, and neither incident type, social engineering or democratic interference, falls within its scope.
@@
AI Act enforcement (C10) The Act’s GPAI provisions, effective August 2025, require adversarial testing, incident reporting, and cybersecurity protections for systemic-risk models (European Union, 2024). The European AI Office’s first Code of Practice (European Commission, 2025) operationalizes the transparency and robustness requirements mapped in Table 6. Early compliance experience confirms our finding that organizations struggle to assemble portable evidence across governance obligations, as no single benchmark suite covers more than half of the rubric at the “strong” level.
6.3. Trade-offs between Performance Measures
The scoring formulation offers a transparent and reproducible summary of coverage breadth across ten criteria. It supports comparison within a release and over time when paired with clear documentation of constructs, metrics, and versioned artifacts. We emphasize that the score is a coverage indicator rather than a ranking of overall safety and must be interpreted with the governance crosswalk and with domain risk.
The Fairness-Privacy Paradox: Reconciling Representativeness with the Right to Anonymity
@@ 6.3. Trade-offs between Performance Measures
A fundamental tension emerging from our synthesis is the operational conflict between the requirement for Representativeness and the rigid Privacy mandates of the General Data Protection Regulation (GDPR) (European Parliament and Council of the European Union, 2016). While mitigating algorithmic bias (Section 4.2.1) necessitates the collection and processing of high-fidelity, demographically diverse datasets to ensure equitable model performance across protected classes, privacy-preserving standards often mandate data minimization and the obfuscation of sensitive attributes. This creates a structural “Fairness-Privacy Paradox”: without access to granular, protected-class identifiers, developers cannot empirically verify statistical parity or conduct intersectional bias audits (Kun, 2015). Consequently, the legal push for anonymity under GDPR can inadvertently mask the very disparities that ResGenAI frameworks seek to eliminate. We posit that resolving this tension requires a transition from absolute data minimization to the adoption of privacy-enhancing technologies (PETs) such as Differential Privacy and Federated Learning that facilitate rigorous fairness auditing while maintaining individual-level anonymity .
Critical Appraisal of Technical Controls: Synthesis and Reflection
After looking closely at how these systems work in the real world, it’s clear there is a significant gap between what we hope these tools can do and how they actually perform when things get messy. The Explanation-Reliability Gap shows us that while we have many ways to explain model decisions (Dwivedi et al., 2023), these often create stories that sound good but don’t actually match what the model is doing under the hood (Strathern, 1997a). We have found that for big models, these explanations are mostly helpful for developers to find bugs, rather than giving regular people or regulators the true transparency they need.
The Bias-Overcorrection Paradox is another hard lesson we’ve learned from cases like the Gemini incident. When we try too hard to fix fairness using simple keywords, we often end up breaking the model’s ability to be factually accurate or culturally aware. This tells us that fairness isn’t just a number to hit (Kun, 2015), but a deep challenge that needs to change based on where the model is being used. Finally, The Shallow Guardrail Problem reminds us that models can look perfectly safe on paper (Zhang et al., 2024) while still being easy to trick. We’ve seen that models which pass every standard toxicity test can still be pushed into strange or harmful behaviour through basic multi-turn conversations (Roose, 2023). The reality is that our current ”guardrails” often just polish the output without fixing the underlying issues, leaving them vulnerable to the automated attacks we see growing today (Samvelyan et al., 2024).
6.4. Analytical Triangulation: Bridging the Evaluation Gap
The persistent safety failures in production systems are not merely isolated technical oversights; they are symptoms of a systemic “Evaluation Gap” where current benchmarks fail to model the adaptive, iterative pressure applied by real-world adversaries. This section triangulates the theoretical risk taxonomy in Section 3 with the empirical deployment failures documented in Section 6.2 through the lens of our comparative assessment (Section 4.2).
Our assessment in Table 4 reveals a stark deficit in evaluative maturity regarding Malicious Actor Realism (Criterion C8), where only 2 out of 21 surveyed benchmarks: Rainbow Teaming and ALERT provide strong coverage for adaptive pressure. This misalignment directly correlates with the high-profile incidents synthesized in Section 6.2, such as the adversarial persona adoption observed in Bing Chat (Roose, 2023) and the scalable data extraction vulnerabilities reported in ChatGPT (Nasr et al., 2023). These failures underscore that prevailing evaluation suites remain largely static and task-local, leaving them unequipped to measure the non-deterministic tactics encountered in the wild.
Furthermore, the representational failures observed in the Gemini overcorrection incident illustrate a critical breakdown in Governance Alignment (C10), where technical alignment efforts are decoupled from the broader sociotechnical context of deployment. By triangulating these findings, we posit that a ResGenAI operating model must transition from one-shot static probes toward continuous, adaptive evaluation frameworks that mirror the malicious-actor realism and distribution shifts inherent in generative ecosystems.
6.5. Limitations of the Study
Like any survey-based study, this work has several limitations. First, the review follows a pre-specified protocol and focuses primarily on the period from November 2022 to December 2025, with selective inclusion of foundational works for definitional clarity. While this post-ChatGPT window captures the most rapid phase of GenAI deployment, it may exclude emerging studies published after the search cutoff or earlier work that did not explicitly use contemporary terminology. Second, our synthesis relies on publicly available papers, dataset cards, benchmark documentation, and governance artifacts. As a result, proprietary evaluations, internal red-teaming practices, or unpublished safety mitigations may be under-represented. Documentation quality also varies across benchmarks and frameworks, which may affect the completeness of criterion-level scoring.
Third, rubric weights are assigned equally by default, and criterion coverage is determined through structured reading of sources followed by discussion and adjudication in cases of disagreement. Although we conducted sensitivity analyses to assess the robustness of equal-weight assumptions, alternative weighting schemes reflecting sector-specific risk priorities could yield different relative rankings. Fourth, the scoring procedure captures breadth of coverage rather than depth or real-world effectiveness. A benchmark receiving a high rubric score indicates multi-risk awareness and governance alignment, but does not guarantee empirical robustness under adversarial deployment conditions. In this sense, the rubric reflects documented capability rather than independently verified performance.
Fifth, the review synthesizes heterogeneous evidence types, including empirical evaluations, conceptual frameworks, and regulatory analyses. This diversity strengthens thematic coverage but precludes formal meta-analysis or quantitative effect-size aggregation. Our conclusions therefore emphasize structural patterns rather than causal claims. Sixth, governance alignment is assessed at the level of plausible evidentiary support rather than legal compliance. Mapping benchmarks to instruments such as the EU AI Act or NIST AI RMF indicates potential contribution to audit evidence, but does not constitute certification or conformity assessment.
Finally, the rapid evolution of multimodal and agentic systems means that new failure modes may emerge after publication. Continuous updates to evaluation suites, governance frameworks, and deployment practices may shift the relative landscape described here. Replication of our screening and scoring is facilitated by the availability of documented sources, explicit inclusion criteria, and the published rubric; however, interpretive judgment remains inherent in structured qualitative synthesis.
6.6. Research directions
The rubric-based analysis in Section 4.2 indicates that commonly used suites provide strong evidence for toxicity and bias, and more limited coverage for privacy, deepfakes, and system-level failures. As a result, many evaluations remain static and task-local, which limits their usefulness for governance where decision makers need evidence that is portable across audits and versions. Future research should target evaluations that model distribution shifts, adaptive adversaries, and multi-agent coordination with external tools.
For practitioners and policy makers, Tables 3 to 8 map benchmarks to risk families and a governance crosswalk provide a starting point for evidence collection. These tables also reveal gaps that are relevant for conformity assessment under emerging regimes. Applied teams can use the matrix to assemble a minimal set of evaluations that cover privacy, robustness, transparency, logging, and post-market monitoring, and then plan additions for multimodal deepfake risk and adaptive red teaming as required by sectoral context.
@@ 6.6. Research directions
We highlight five priorities. First, we need to extend benchmarks to capture system-level failures in tool-using and multi-agent settings, especially around retrieval, planning, and actuation (e.g., API calls, code execution, or robotic steps). Second, fold privacy and provenance checks into routine testing, including membership-inference and training-data extraction probes, plus audits of logging sufficiency. Third, create deepfake risk assessments for multimodal models that cover detection, content provenance, and robustness to evasion. Fourth, shift from static to adaptive evaluation with undisclosed items, targeted red teaming, and tests under distribution shift. Fifth, report sustainability alongside accuracy and robustness energy, emissions, and compute-normalized efficiency to enable compute-proportional evaluation and responsible deployment. These priorities directly address the risk-surface gaps surfaced in this survey.
Teams putting ResGenAI into practice can follow a simple loop. First, scope risks by mapping planned features to common risk families (e.g., bias, toxicity, privacy, and security). Next, select baselines from our matrix so that each release is covered by minimum tests across those areas. Then, add context: turn on logging, enforce dataset/model versioning, and keep lightweight documentation so decisions and results remain traceable. Before launch and later during post-market monitorin, there is need to exercise adaptivity with red teaming and a few undisclosed checks to catch regressions. Finally, track a small set of KPIs across the lifecycle: calibration, error rates in high-stakes flows, audit cadence, exclusivity indicators, and energy/emissions. This routine keeps evaluation aligned with governance expectations and makes evidence easier to reuse when audits occur.
6.7. Systemic Vulnerabilities and Blind Spots in ResGenAI
While GenAI capabilities have accelerated, structural blind spots persist across technical and sociotechnical domains (Kenny et al., 2021). This review identifies several systemic gaps that currently hinder universal assurance:
  • RLHF and Alignment Transparency: Alignment advances, such as RLHF and Constitutional AI, have shifted emphasis from raw capability to controllability. However, the documentation of these alignment pathways is often inconsistent, which complicates the auditing of model refusal behaviour and evidence portability across versions (of Standards and, NIST, NIST).
  • Opaque Proprietary Datasets: Responsible stewardship demands adherence to consent and privacy regimes. However, many frontier models rely on web-scale datasets where provenance, licensing, and curation criteria remain undocumented. Truly representative datasets are often infeasible due to data scarcity or legal limits, leading to persistent data leakage and memorization risks (Nasr et al., 2023).
  • Non-Stationary Safety: Finally, safety is inherently non-stationary; model updates, distribution shifts, and evolving adversarial tactics induce drift. Static benchmarks rarely capture these dynamic failure modes , necessitating a shift toward the continuous monitoring and adaptive evaluation frameworks discussed in Section Section 4.2 .

### v7

@@ 6. Discussion
Post-GPT scholarship concentrates on five areas: safety and benchmarking, fairness and governance, explainability and transparency, privacy and security, and sectoral applications. Bias and toxicity receive dense coverage; deepfakes, sustainability, and system-level failure modes receive comparatively less. This pattern aligns with the growing prominence of multimodal and agentic systems in 2022–2025, which heightens demand for evaluations capturing tool use and context.
@@
To fully answer Who is Responsible?, our synthesis posits that users must transition from passive consumers to systemic guardrails. The current imbalance—74% of organizations use GenAI but only 26% have a formal strategy—routinely leaves end-users inheriting unmitigated risks such as PII leakage and hallucination. Triangulating the sectoral failures in Section 5 with the digital-literacy gaps in Section 3.2, we propose a Symmetric Responsibility Model: developers are accountable for Transparency and Alignment; users are accountable for Operational Integrity and Dissemination. Digital literacy is therefore not merely a skill but a core technical control, assessable alongside model-side benchmarks (Section 4.2). Without user-centric governance, even perfectly aligned models remain vulnerable to last-mile failures of misinterpretation or malicious recontextualization. Appendix Table 5 operationalizes this model by mapping failure modes to the reciprocal duties of each stakeholder.
@@
Industry Safety Artifacts and Advanced Interpretability. While academic benchmarks offer standardized metrics, the operational reality of ResGenAI is increasingly shaped by safety documentation from frontier labs and by emerging mechanistic interpretability techniques (Figure 10). OpenAI system cards for GPT-4 and GPT-4o provide granular audits of sycophancy and PII leakage; Anthropic’s Constitutional AI automates alignment via RLAIF with tiered safeguards; Meta Llama 3 introduces standardized multilingual red-teaming protocols; and Google DeepMind reports detail bias–accuracy trade-offs and overcorrection mechanisms. These artifacts map directly to our C1–C10 rubric and function as de facto governance-facing benchmarks. Complementing this, mechanistic interpretability techniques such as Sparse Autoencoders (SAEs) isolate circuits responsible for deceptive or biased reasoning, enabling targeted model steering during inference—addressing the black-box challenge through empirical verification of internal safety alignment before deployment.
@@
Figure 10. Integration of industry safety documentation and advanced mechanistic interpretability for governance-facing evidence.
@@
ResGenAI Sustainability via Lifecycle Assessment. Sustainability must transition from qualitative discourse to rigorous quantification grounded in Lifecycle Assessment (LCA) frameworks, as integrated into Figure 7 and synthesized in Appendix Figure 2. Following the KPIs in Section 4.2.4, we advocate reporting energy consumption and carbon-equivalent emissions normalized by throughput (e.g., per 1K tokens) across pretraining, fine-tuning, and inference. Training large-scale generative models produces environmental footprints comparable to the lifetime emissions of multiple automobiles, underscoring the need for transparent tracking tools and hardware-aware reporting.
Lessons from Industry Deployment Incidents. Real-world incidents expose risk-coverage gaps that static benchmarks miss.
@@
Alignment and safety failures (C2, C3, C8). Microsoft Bing Chat’s GPT-4 integration produced adversarial persona adoption and emotional manipulation through multi-turn prompting (Roose, 2023)—an adaptive adversary pattern that only 2 of 21 benchmarks address. OpenAI’s GPT-4o system card documented sycophantic and manipulative voice-mode outputs beyond the reach of toxicity classifiers. Google’s Gemini image generator was suspended in February 2024 after overcorrection in bias mitigation produced historically inaccurate outputs, illustrating how fairness interventions can themselves become failure modes.
@@
Hallucination in deployed systems (C4, C9). In Mata v. Avianca (2023), attorneys submitted six fabricated ChatGPT-generated case citations, resulting in judicial sanctions.333caselaw.findlaw.com Only 7 of 21 benchmarks address misinformation with documented constructs, confirming that the high-stakes error rate $E_{h}$ (Section 4.2.4) must be operationalized beyond laboratory settings.
@@
Privacy and data leakage (C6). Samsung banned internal ChatGPT use in 2023 after employees leaked proprietary source code (Samsung Electronics, 2023)—a usage-path vector no current privacy benchmark assesses. Nasr et al. (Nasr et al., 2023) demonstrated scalable extraction of memorized PII from production ChatGPT via divergence attacks, exceeding the scope of PrivLM-Bench and PrivacyLens.
@@
Deepfake and media integrity (C5). A Hong Kong finance worker transferred \$25M after a deepfake video call impersonating executives (Chen, 2024); AI-generated robocalls impersonating President Biden targeted the 2024 New Hampshire primary (New Hampshire Department of Justice, 2024). Both incidents fall outside the evaluation scope of all but one benchmark (SHIELD).
@@
AI Act enforcement (C10). The EU AI Act’s GPAI provisions (effective August 2025) require adversarial testing, incident reporting, and cybersecurity protections for systemic-risk models (European Union, 2024). The European AI Office’s first Code of Practice (European Commission, 2025) operationalizes the transparency and robustness requirements mapped in Table 5. Early compliance analysis confirms that no single benchmark suite covers more than half the rubric at the “strong” level.
Trade-offs between Performance Measures
@@
The Fairness-Privacy Paradox. Mitigating algorithmic bias requires demographically granular data to verify equitable performance, yet privacy standards such as GDPR (European Parliament and Council of the European Union, 2016) mandate data minimization and obfuscation of sensitive attributes. This tension means that legal anonymity requirements can inadvertently mask the disparities ResGenAI frameworks seek to eliminate: without protected-class identifiers, developers cannot empirically verify statistical parity or conduct intersectional audits (Kun, 2015). Resolution requires moving beyond absolute minimization toward privacy-enhancing technologies (PETs), such as differential privacy, federated learning, secure multiparty computation, which enable rigorous fairness auditing while preserving individual-level anonymity.
Critical Appraisal of Technical Controls. Three recurring gaps emerge from deployment evidence. The Explanation–Reliability Gap: post-hoc explanations (Dwivedi et al., 2023) frequently construct plausible narratives that do not reflect internal model computations (Strathern, 1997), limiting their value beyond developer debugging. The Bias–Overcorrection Paradox: as the Gemini incident illustrates, aggressively enforcing keyword-level fairness targets can degrade factual accuracy and cultural awareness, confirming that fairness is a context-dependent sociotechnical property rather than a fixed metric (Kun, 2015). The Shallow Guardrail Problem: models that pass standard toxicity benchmarks (Zhang et al., 2024) remain exploitable via basic multi-turn conversations (Roose, 2023) and automated adversarial attacks (Samvelyan et al., 2024), indicating that surface-level output filtering does not address underlying alignment deficits.
Systemic Vulnerabilities, Blind Spots, and Study Limitations
Three structural blind spots persist across the field. (1) RLHF and Alignment Transparency: Alignment advances such as RLHF and Constitutional AI have shifted emphasis toward controllability, but inconsistent documentation of alignment pathways complicates auditing of refusal behavior and evidence portability across model versions. (2) Opaque Proprietary Datasets: Many frontier models rely on web-scale datasets where provenance, licensing, and curation criteria remain undocumented, sustaining persistent data-leakage and memorization risks. (3) Non-Stationary Safety: Model updates, distribution shifts, and evolving adversarial tactics induce drift that static benchmarks rarely capture, necessitating continuous monitoring and adaptive evaluation.
These field-level gaps are compounded by limitations of the present study. Our protocol covers 2022–2025 with selective inclusion of foundational works; studies published after the search cutoff or using non-contemporary terminology may be excluded. Synthesis relies on publicly available sources, so proprietary evaluations and unpublished safety mitigations may be under-represented. Rubric weights are equal by default; sector-specific schemes could yield different rankings. Scores capture breadth of coverage rather than depth or real-world effectiveness. Evidence types are heterogeneous—empirical evaluations, conceptual frameworks, regulatory analyses—precluding formal meta-analysis; conclusions therefore emphasize structural patterns rather than causal claims. Finally, the rapid evolution of multimodal and agentic systems means new failure modes may emerge after publication.
Research Directions
@@
We highlight five priorities. First, extend benchmarks to capture system-level failures in tool-using and multi-agent settings (retrieval, planning, actuation). Second, fold privacy and provenance checks—membership-inference, training-data extraction probes, and logging audits—into routine testing. Third, create deepfake risk assessments for multimodal models covering detection, content provenance, and evasion robustness. Fourth, shift from static to adaptive evaluation with undisclosed items, targeted red teaming, and distribution-shift testing. Fifth, report sustainability (energy, emissions, compute-normalized efficiency) alongside accuracy and robustness to enable compute-proportional evaluation and responsible deployment. These priorities directly address the risk-surface gaps surfaced in this survey.